Highsmith Rental Company purchased an apartment building early in 2021. There are 20 apartments in the building and each is furnished with major kitchen appliances. The company has decided to use the group depreciation method for the appliances. The following data are available:

Appliance Cost Residual Value Service Life
(in Years)
Stoves $ 39,000 $ 3,000 6
Refrigerators 34,000 4,000 5
Dishwashers 32,000 3,000 4

In 2024, two new refrigerators costing $3,900 were purchased for cash. In that same year, the old refrigerators, which originally cost $3,900, were sold for $1,400.

Required:
1. Calculate the group depreciation rate, group life, and depreciation for 2021.
2. Prepare the journal entries to record the purchase of the new refrigerators and the sale of the old refrigerators.
Answer #1

Ans:

Computation of Depreciation
Asset Cost(a) Residual Value (b) Depreciable Base Estimated Life Annual Depreciation(SLM)
(c=a-b) (years) (d) (c/d)
Stoves $ 39,000.00 $ 3,000.00 $ 36,000.00 6 $ 6,000.00
Refrigerators $ 34,000.00 $ 4,000.00 $ 30,000.00 5 $ 6,000.00
Dishwashers $ 32,000.00 $ 3,000.00 $ 29,000.00 4 $ 7,250.00
Totals $ 105,000.00 $ 10,000.00 $ 95,000.00 $ 19,250.00

Computation of Group Depreciation Rate, Group Life

Group Depreciation Rate= $19250/105000= 18.33%

Group Asset Life= 95000/19250= 4.94 Rounded Off

NO Event General Journal Debit Credit
1 1 Refrigerators $ 3,900.00
Cash $ 3,900.00
2 2 Cash $ 1,400.00
Accumulated depreciation $ 2,500.00
Refrigerators $ 3,900.00